# Prague, Czech

### v3

@@ 6 Discussion
Our RL framework for precision selection requires proper state modeling of the data. For instance, we use features such as the approximate condition number and matrix norm in our state modeling for the linear solve problem. Selecting or modeling the states often requires expert knowledge to balance reasonable computational complexity with sufficient information for effective RL training.
However, this limitation, on the other hand, can also be viewed as advantages. In this work, the proposed framework has shown efficient precision selection even when prior knowledge about the required precision for critical computational steps is lacking. Although prior studies have explored mixed-precision usage in iterative refinement methods based on condition numbers and matrix sizes, our framework offers two unique benefits. First, it can quickly determine the required precision for unseen linear systems. Second, it allows us to identify the required precision for selected matrix features in a black-box setting. For example, by selecting certain features in our state space, we can examine whether these features are key factors (salient or not) that determine the reduced mixed precision. When the salient features are unknown or when their influence on the required precision level is unclear, our framework becomes particularly advantageous.
@@ 7 Conclusion and Future Work
We propose a reinforcement learning framework for automated precision selection for linear solvers by modeling a contextual bandit problem, with a case study in iterative refinement. This framework enables adaptive precision selection for customized computational steps, thereby optimizing the trade-off between numerical accuracy and computational efficiency. Our empirical results show that RL, leveraging a discretized state space, effectively adapts precision choices to the numerical properties of each linear system and achieves a good compromise between algorithm performance (relative error and converged iterations) and the reduced-precision portion of the linear systems under varying conditions.

### v4

@@ 6 Discussion
Unlike supervised learning problem, such as classification, our contextual bandit RL framework is more well-suited for practical solve since it does not require ground-truth labels, which is of more practical to numerical analysis problems. However, our RL framework for precision selection requires proper state modeling of the data. For instance, we use features such as the approximate condition number and matrix norm in our state modeling for the linear solve problem. Selecting or modeling the states often requires expert knowledge to balance reasonable computational complexity with sufficient information for effective RL training.
However, this limitation, on the other hand, can also be viewed as an advantage. In this work, the proposed framework has shown efficient precision selection even when prior knowledge about the required precision for critical computational steps is lacking. Although prior studies have explored mixed-precision usage in iterative refinement methods based on condition numbers and matrix sizes, our framework offers two unique benefits. First, it can quickly determine the required precision for unseen linear systems. Second, it allows us to identify the required precision for selected matrix features in a black-box setting. For example, by selecting certain features in our state space, we can examine whether these features are key factors (salient or not) that determine the reduced mixed precision. When the salient features are unknown or when their influence on the required precision level is unclear, our framework becomes particularly advantageous.
@@ 7 Conclusion and Future Work
We present a reinforcement learning framework for automated precision selection in linear solvers, casting the problem as a contextual bandit that balances convergence across iterations and low precision usage. Using iterative refinement as a case study, the framework dynamically adjusts precision at different computation steps to better balance numerical accuracy and efficiency. Our empirical results show that RL, leveraging a discretized state space, effectively adapts precision choices to the numerical properties of each linear system and achieves a good compromise between algorithm performance (relative error and converged iterations) and the reduced-precision portion of the linear systems under varying conditions.